### v2

# Dark Matter (H)eats Young Planets

Abstract
We study the effect of dark matter annihilation on the formation of Jovian planets. We show that dark matter heat injections can slow or halt Kelvin-Helmholtz contraction, preventing the accretion of hydrogen and helium onto the solid core. The existence of Jupiter in our solar system can therefore be used to infer constraints on dark matter with relatively strong interaction cross sections. We derive novel constraints on the cross section for both spin-dependent and spin-independent dark matter. We highlight the possibility of a positive detection using future observations by JWST, which could reveal strongly varying planet morpholoiges close to our Galactic Center.
preprint: IPPP/23/43preprint: LTH-1349
Our technical capabilities to detect planets outside our solar system have developed dramatically. From the first indirect evidence for an exoplanet Mayor and Queloz (1995), to direct imaging Chauvin et al. (2005) and precision atmospheric observations Martins et al. (2015), we have now 5.5× $5.5\times\,10^{3}$ detected exoplanets, and $\mathcal{O}(10^{4})$ candidates awaiting confirmation cat (2020). Especially with the onset of high performance infra-red astronomy and the JWST Brande et al. (2019), and Roman Johnson et al. (2020) space missions low temperature targets, such as exoplanets can be studied across the galaxy, all the way to the Galactic Center (GC). This development opens up a new window to our galaxy, and allows us to study exoplanets in various astrophysical environments.
Dark Matter (DM) as the dominant mass source in our galaxy is well mapped by stellar kinematics, and despite uncertainties displays a density profile that raises towards the GC Benito et al. (2021). In the search for this ellusive substance a number of studies have considered DM capture in celestial object ranging from the Earth and the Sun Batell et al. (2010); Pospelov et al. (2008); Pospelov and Ritz (2009); Rothstein et al. (2009); Chen et al. (2009); Schuster et al. (2010a, b); Bell and Petraki (2011); Feng et al. (2016a); Kouvaris and Tinyakov (2010a); Feng et al. (2016b); Allahverdi et al. (2017); Leane et al. (2017); Arina et al. (2017); Albert et al. (2018a, b); Nisa et al. (2019); Niblaeus et al. (2019); Cuoco et al. (2020); Serini et al. (2020); Acevedo et al. (2021); Mazziotta et al. (2020); Bell et al. (2021), Jupiter Batell et al. (2010); Leane and Linden (2021); Li and Fan (2022); French and Sher (2022); Ray (2023), Brown Dwarfs Leane and Smirnov (2021); Leane et al. (2021), Uranus Mitra (2004), White Dwarfs Garani et al. (2023), Neutron Stars Goldman and Nussinov (1989); Gould et al. (1990); Kouvaris (2008); Bertone and Fairbairn (2008); de Lavallaz and Fairbairn (2010); Kouvaris and Tinyakov (2010b); McDermott et al. (2012); Kouvaris and Tinyakov (2011a); Guver et al. (2014); Bramante et al. (2013); Bell et al. (2013); Bramante et al. (2014); Bertoni et al. (2013); Kouvaris and Tinyakov (2011b); McCullough and Fairbairn (2010); Angeles Perez-Garcia and Silk (2015); Bramante (2015); Graham et al. (2015); Cermeno et al. (2016); Graham et al. (2018); Acevedo and Bramante (2019); Janish et al. (2019); Krall and Reece (2018); Baryakhtar et al. (2017); Raj et al. (2018); Bell et al. (2018); Chen and Lin (2018); Dasgupta et al. (2019); Hamaguchi et al. (2019); Camargo et al. (2019); Bell et al. (2019); Acevedo et al. (2020); Joglekar et al. (2019, 2020); Bell et al. (2020); Garani et al. (2020); Leane et al. (2021); Bramante et al. (2022a); Coffey et al. (2022), and other stars Freese et al. (2009); Taoso et al. (2008); Ilie et al. (2020a, b); Lopes and Lopes (2021); Ellis (2021); Croon and Sakstein (2023), as well as recently Exoplanets Leane and Smirnov (2021); Bramante et al. (2022b) (for a recent review, see Bramante and Raj (2023)). Many of the above studies rely on an energy injection from DM particles that annihilate after being captured by the celestial object. This is expected in models where DM has been thermally produced in the early universe.
Jovian exoplanets have been identified as ideal candidates for DM capture in Ref. Leane and Smirnov (2021). The advantage is their relatively large surface gravity, combined with rather low core temperatures, leading to DM with low masses to be retained in the object. Their hydrogen-rich composition makes them kinematically well suited for capturing DM particles around the GeV scale, and especially particles with spin-dependent nuclear interactions. The differential measurement of Jovian planet and Brown Dwarf temperatures close to the GC has been identified as a promising tool to detect heating from annihilating DM  Leane and Smirnov (2021).
In this work we propose a more radical possibility. By investigating the formation process of Jovian planets we argue that an additional heat source associated with DM annihilation could halt the formation process at various stages, depending on the DM parameters, such as mass, cross section, and ambient number density. Thus, we would expect to have varying properties of Jovian planets across the galaxy, or their total absence under certain circumstances.
Figure 1: Benchmark for spin-dependent dark matter: $m_{\rm DM}=10$ GeV, $\sigma_{p}=4\times 10^{-28}\,\rm cm^{2}$. The red line shows the dark matter capture fraction at the time when the planet is $10\,\rm M_{\oplus}$ and the black continuous and dashed lines show the fractions needed to halt formation through gas capture. This fraction depends on the temperature of the surrounding gas, and thus on the proximity to the host star.
In Fig. 1 we show the capture efficiency as a function of distance from the GC assuming an NFW profile and a Maxwellian velocity dispersion. As shown here, at a given distance from the host star (corresponding to a given external temperature $T_{\rm ext}$), planet formation can be halted in the GC while it does occur some distance away from it. This would imply that towards the GC, Jovian planets appear to occur further closer to the host stars, modulo migration after formation. This observation provides a new and radically different search for a DM heating signal within our galaxy. This distinct phenomenology can be targeted by future observations of exoplanets leveraging the advances in the development of infrared telescopes.
Finally, the mere existence of Jupiter in our solar system allows us to set new constraints on the parameter space of DM. We find that future observations of exoplanets near the GC, enabled by the JWST and future infrared telescopes will further test yet unconstrained parameter space of DM candidates across a wide range of masses and interaction cross sections.
Planet formation
We focus on giant, Jovian planets, and the formation process called core-accretion gas-capture Hubickyj et al. (2005), which proceeds in two stages. First, a solid core made of heavier elements, such as ice and silica forms within a fraction of a Myr. These planets form beyond the frost line of a solar system, where icy compounds are cool enough to remain solid, allowing cores to become heavy enough to start to accrete hydrogen and helium. Planets formed closer to their sun, such as Earth, on the inside the frost line have cores dominated by iron and silicates, which are too rare to lead to large cores that would accrete gas.
Gas capture does not begin until the core accretion exceeds about $10M_{\oplus}$ ($M_{\oplus}=6.0\times 10^{24}$ kg), when the growing core raises the escape velocity above the thermal velocity of the nebular gas. The proto-planet then accumulates a gas envelope at an increasing rate, reflecting its increasing mass. The gas can quickly fill up the Bondi sphere of the planet, which subsequently needs to cool via the Kelvin-Helmholtz mechanism in order to accrete further Pollack et al. (1996); Adams and Batygin (2022), until eventually the protoplanet reaches a threshold mass at which runaway accretion takes place. Therefore, during the gas accretion period before runaway (which takes place on the scale of a Myr  D’Angelo et al. (2021)) the system remains in a quasi-equilibirum state where gas accretion can only proceed once the excess heat that is released from the gravitational binding energy has been radiated away. It is then not surprising that an additional heat source can disrupt this fragile equilibrium and lead to drastically different outcomes. Note in particular that the hydrogen accretion stage is unique, as hydrogen atoms have the lowest gravitational binding energy and are thus most susceptible to evaporation from the forming planet.
Disrupting Jovian Formation
We assume that DM is produced in a scenario resulting in equal abundances of DM particles and antiparticles. In the presence of elastic scattering interactions with standard model (SM) particles, DM can scatter in forming planets, lose its kinetic energy, and become gravitationally captured. Once its number density is large enough for particle-anti-particle annihilation to become efficient, it can become an additional substantial heat source in the system. Thus, it may slow or halt Kelvin-Helmholtz contraction, in turn preventing the accretion of gas and the formation of Jovian planets. The effect becomes relevant once the energy injection due to DM annihilation is comparable to the power generated by the Kelvin-Helmholtz mechanism.
Contraction is set by the net energy loss, which is modified in the presence of DM annihilation,
$\begin{split}\dot{Q}-L_{\rm DM}(R)&=-\frac{dU}{dt}\\ 4\pi R^{2}\times\sigma_{\rm SB}T^{4}-L_{\rm DM}(R)&=c_{1}\frac{GM^{2}}{R^{2}}\frac{dR}{dt}\end{split}$ (1)
where $\dot{Q}$ is the radiative heat loss, $dU/dt$ the change in gravitational potential energy, $L_{\rm DM}(R)$ is the $R$ dependent heat injection due to DM annihilation, which we call DM luminosity (note that this is not necessarily directly observable as luminosity of the protoplanet), and $1/2\leq c_{1}\leq\infty$ is a factor which depends on the internal mass distribution (for a uniform sphere, $c_{1}=3/5$).
Initially, dark matter heat injections will raise the temperature of the protoplanet, increasing its radiative heat loss. However, above a certain temperature the mixture of hydrogen and helium gas acquires enough thermal energy such that particles near its surface reach the escape energy and become unbound.
This process will be in competition with the mass accretion rate, which we take from numerical simulations Lissauer et al. (2009). To quantify the particle loss rate we consider the generalisation of the Jeans escape rate $\Gamma^{0}_{\rm J}(T)$ Jeans (2009) by including the effect of a chemical potential difference
$\Gamma_{\rm J}(T)=\Gamma^{0}_{\rm J}(T)\exp\left(-\frac{\Delta\mu}{N\,T}\right)$
$=\frac{nv}{2\sqrt{\pi}}\left(1+\frac{v_{\rm esc}^{2}}{v^{2}}\right)\exp\left(-\frac{E_{\rm esc}}{T}\right),$ (2)
where $n$ is the number density of gas particles, $v_{\rm esc}$ the surface escape velocity, $v=\sqrt{2T/m}$, and the escape energy that takes into account the difference in chemical potentials
$E_{\rm esc}=\frac{G_{N}Mm}{r}+\frac{\Delta\mu}{N}=\frac{m\,v_{\rm esc}^{2}}{2}+\frac{\Delta\mu}{N}\,.$ (3)
Here, $G_{N}$ is the Newton constant, $M$ the mass of the forming planet, $m$ the mass of the evaporating particle, and $\Delta\mu/N$ is the difference in chemical potentials per particle. It is intriguing that the chemical potential is relevant in this case, reflecting the special situation of evaporation of particles from the planetary envelope into the surrounding gas of the protoplaneary disk, which has a comparable particle number density and temperature.
Given that $\mu_{1}$ is the chemical potential in the envelope and $\mu_{2}$ that of the protoplaneraty disk, the difference in the chemical potentials $\Delta\mu$ is given by
$\mu_{1}-\mu_{2}=U_{1}-U_{1}+P_{1}V_{1}-P_{2}V_{2}+T_{2}S_{2}-T_{1}S_{1}$ (4)
where we assumed that the gravitational energies of the particles are of the same order. We further assume that the particles behave as ideal gas, and thus the internal energy is $U_{i}=C_{V}N_{i}T_{i}$, $P_{i}V_{i}=N_{i}T_{i}$, and the entropy is given by the Sackur-Tetrode formula $S_{i}=N_{i}\kappa_{i}$, with
$\kappa_{i}=\left(\frac{5}{2}+\log\left[\frac{V_{i}}{N_{i}}\left(\frac{4\pi mU_{i}}{3N_{i}}\right)^{3/2}\right]\right)\,.$ (5)
Furthermore, we assume that the gases are monoatomic, and hence the heat capacity $C_{V}\approx 3/2$.
We can therefore express the chemical potential difference per envelope particle as
$\frac{\Delta\mu}{N_{1}}=C_{V}\left(T_{1}-\frac{N_{2}}{N_{1}}T_{2}\right)+T_{1}-\frac{N_{2}}{N_{1}}T_{2}+\bar{\kappa}\left(T_{2}\frac{N_{2}}{N_{1}}-T_{1}\right)$
$\approx\left(1+C_{V}-\bar{\kappa}\right)\left(T_{1}-T_{2}\right)=\frac{3}{2}\log\left[\frac{4\pi m\bar{u}}{3\bar{n}^{2/3}}\right]\left(T_{2}-T_{1}\right)\,.$ (6)
Where we have used the assumption that $N_{1}\approx N_{2}$, and as the logarithm is a slowly varying function, the argument can be replaced by the average value of the internal energy per particle $\bar{u}=\bar{U}/\bar{N}\approx C_{V}\bar{T}$. Thus, we observe that in the case that the surrounding gas is hotter than the envelope an additional suppression slows particle evaporation down.
Now we compare the escape rate with the rate of growth of the planet in the absence of dark matter heat $\dot{M}$, which we take from simulation results presented in Lissauer et al. (2009). Equating those rates
$A_{\rm planet}\,\Gamma_{\rm J}(T)|_{T=T_{\rm max}}=\mathcal{O}(1)\,\dot{M}$ (7)
where $A_{\rm planet}$ is the surface from which evaporation takes place, we find that the result is insensitive to the value of the numerical $\mathcal{O}(1)$ factor in the equation, since it can not compete with the exponential in the escape rate (see also D’Angelo et al. (2021), which is largely consistent but presented slower accretion, implying smaller DM injections are needed to match it).
For an ambient temperature of $115$ K of the surrounding gas Lissauer et al. (2009), this treatment gives us hydrogen escape temperatures of $T_{\rm max}\sim 85$ K for the envelope. For the core the assumption above has to be modified given that the particle number density in the core is much higher than in the surrounding gas. We get $T_{\rm max}\sim 10^{3}$ K for the core, a result that is not significantly affected by the ambient gas properties for the systems we study.
Thus, the maximum heat loss happens just below $T_{\rm max}$ and we estimate that DM heat injection halts further gas capture if
$\begin{split}L_{\rm DM}&\geq 4\pi R^{2}\sigma T_{\rm max}^{4}\\ &\geq 4\times 10^{-4}\times\left(\frac{R}{10^{3}R_{p}}\right)^{2}\,\left(\frac{T_{\rm max}}{80\rm K}\right)^{4}\rm\,L_{\odot}\end{split}$ (8)
where $R_{p}=7\times 10^{5}$ km is the radius of Jupiter (the protoplenatary envelope has radius $\sim 10^{3}R_{p}$), and where $L_{\odot}=2\times 10^{36}\,\rm GeV\,s^{-1}$ is the luminosity of the Sun. To investigate whether Eq. (8) is satisfied in the evolution of the planet, we use the simulation results presented in D’Angelo et al. (2021). We will now examine how the DM luminosity depends on the properties of the capturing object, and the DM particle model.
Dark matter energy injection
To obtain the DM luminosity, we assume DM annihilation equilibrium – motivated a posteriori below – i.e. that each incoming DM particle that is captured contributes its rest mass immediately to the energy budget. We will discuss why this treatment is justified in the following subsection. We thus can write
$L_{\rm DM}=m_{\rm DM}C_{\rm cap}=m_{\rm DM}f_{\rm cap}\Phi\pi R^{2}$ (9)
where $C_{\rm cap}$ is the capture rate, and $f_{\rm cap}$ the captured fraction of the DM flux $\Phi$ passing through the object, given by
$\Phi=\,v_{\rm DM}\sqrt{\frac{8}{3\pi}}\frac{\rho_{\rm DM}}{m_{\rm DM}}\left(1+\frac{3}{2}\frac{v_{\rm esc}^{2}}{v_{\rm DM}^{2}}\right)\,.$ (10)
Then, we have
$\begin{split}L_{\rm DM}\sim&7\times 10^{-4}f_{\rm cap}\left(\frac{R}{10^{3}R_{p}}\right)^{2}\\ &\times\left(\frac{v_{\rm DM}}{270\rm km\,s^{-1}}\right)\left(\frac{\rho_{\rm DM}}{0.42\,\rm GeV\,cm^{-3}}\right)\,L_{\odot}\end{split}$ (11)
which only depends on the dark matter mass through $f_{\rm cap}$. It is seen that this may indeed exceed (8).
Dark matter capture
Given a certain DM mass and velocity, the DM particle needs a certain number of scatterings to drop below the escape velocity of a given object. This will in general also depend on the efficiency of the energy loss, determined by the mass ratio of the DM particle and the target nucleus – DM is not sufficiently slowed down when this mass ratio is too large. Furthermore, in the case of DM particles with masses similar to or below the target nucleus mass, DM particles might be reflected off the object before capture can occur, as discussed in detail in Ref. Leane and Smirnov (2023a, b). To compute the fraction of the DM particles that are captured in a given object we use the Asteria Leane and Smirnov (2023b, a) package, which takes all the above mentioned phenomena into account.
In this work we focus on DM interaction with nuclei and distinguish two broad DM model classes. The first is a model where the DM particle scatters with SM particles independently of their spin. Given an input DM nucleon cross section $\sigma_{\chi N}$, this leads to the spin-independent DM-nucleus cross section with an atomic number $A$ nucleus
$\sigma_{\rm SI}=\sigma_{\chi N}^{\rm SI}A^{2}\,\left(\frac{\mu(m_{\chi},m_{A})}{\mu(m_{\chi},m_{N})}\right)^{2}\,,$ (12)
where $\mu(m_{\chi},m_{\rm SM})$ is the reduced mass of a DM and a SM particle of mass $m_{\rm SM}$. Note that it has been pointed out that at cross sections $\sigma_{\chi N}>10^{-26}\text{cm}^{2}$ the scattering is not described by interactions of point like particles, and the coherent scattering picture breaks down Digman et al. (2019); Cappiello et al. (2021). Rather, at cross sections larger than this threshold the interaction can be assumed as a total geometric scattering cross section with the entire nucleus without an implied scaling with atomic number.
The second case of interactions we consider is a spin-dependent scattering cross section
$\sigma_{SD}=\sigma_{\chi N}^{\rm SD}\,\frac{4\left(J_{A}+1\right)}{3J_{A}}\large[a_{p}\langle S_{p}\rangle+a_{n}\langle S_{n}\rangle\large]^{2}\,\left(\frac{\mu(m_{\chi},m_{A})}{\mu(m_{\chi},m_{N})}\right)^{2}\,,$ (13)
where $J_{A}$ is the nuclear spin of the target, $\langle S_{p}\rangle$ the average proton and $\langle S_{n}\rangle$ the average neutron spin of the nucleus. The $a_{p}$, and $a_{n}$ are the DM model parameters indicating the coupling strength to protons and neutrons respectively. For example in the case of $a_{p}=1$, and $a_{n}=0$, the hydrogen scattering cross section is equal to the DM scattering cross section on protons $\sigma_{SD}=\sigma_{p}$, which we will use as a benchmark model.
To compute the capture of DM in the protoplanet we introduce a simplifying effective framework and use the Asteria Leane and Smirnov (2023b, a) package to compute capture in two regimes depending on where in the object we expect the DM to be captured. In the first regime the DM capture takes place on the scale of the entire object and the average chemical composition of the protoplanet is taken as an input for Asteria, while in the other regime the capture takes place in the envelope. In the second case only the chemical composition of the envelope is used, but the total mass of the protoplanet determines the escape velocity. To distinguish the regimes we define the effective capture length $\lambda_{\rm cap}$ and compare it to the size of the protoplanet, which is dominated by the size of the envelope.
$\lambda_{\rm cap}=\lambda_{\rm mfp}\,N_{\rm cap}\text{ with }\lambda_{\rm mfp}=\left(\sigma_{\rm SI/SD}\,n_{\rm SM}\right)^{-1}\,,$ (14)
$N_{\rm cap}=\frac{\log\left(v_{\text{esc}}^{2}/(v_{\chi}^{2}+v_{\text{esc}}^{2})\right)}{\log\left(1-2\mu(1-\mu)^{-2}\right)}\,,$
where $N_{\rm cap}$ is the number of scatterings required to capture a DM particle Mack et al. (2007), $v_{\text{esc}}$ the protoplanets’ escape velocity, $v_{\chi}$ the DM impact velocity, and $\mu=m_{\rm DM}/m_{\rm SM}$ is the mass ratio between the DM and the average SM particle in the protoplanet. Note that this estimate provides an upper bound on the capture length of DM as a straight-line trajectory is assumed.
To determine the chemical composition we model the protoplanet in a simplified manner assuming that it consists of an icy core are an outer gas envelope, both are assumed to have a homogeneous average density. We use the mass and radius from Ref. Lissauer et al. (2009) as time dependent input quantities. The outer envelope has a large radius and a low density. We model the envelope to consist of 75% hydrogen and 25% helium. The core is assumed to have a $58/42$ ice ($\rm H_{2}O$) to silica ($\rm SiO_{2}$) mass ratio D’Angelo et al. (2021), which makes Oxygen the most abundant element. This implies that in the regime where the core contributes to the capture, optimal capture happens for slightly heavier DM candidates, due to more efficient energy transfer between the SM and DM particles of similar mass.
For the spin-dependent calculation, we need to estimate the mass fraction which is relevant for spin-dependent scattering. Thus we take the isotope fractions and properties of the elements in Tab. 1.
Table 1: Isotope fractions, nuclear spins, average proton, and neutron spins of the dominant isotopes considered.
[…]
J 5/2 1/2 1/2
$\langle S_{p}\rangle$0.036 0.054 0.5
$\langle S_{n}\rangle$ 0.508 0.204 0
Lastly, we note that light DM candidates can be reflected, such that the capture efficiency remains below one even at large cross sections Leane and Smirnov (2023a). On the other hand, in models with a long range force mediator, the escape velocity in the medium can be increased, which in turn suppressed the reflection probability Acevedo et al. (2023). We discuss the implications for the cross section sensitivity estimate in the following section.
Dark matter annihilation
Neglecting DM evaporation from the celestial body, the Boltzmann equation governing the accumulation of DM is given by
$\frac{dN}{dt}=C_{\rm cap}-C_{\rm ann}N_{\chi}^{2}\,,$ (15)
where $C_{\rm cap}$ is the total capture rate of DM particles per unit time in the celestial object, and $C_{\rm ann}=\langle\sigma_{\rm ann}v\rangle/V_{\rm eff}$ the DM annihilation rate per effective volume, which is given by $1/V_{\rm eff}=\int_{V}n_{\chi}^{2}/\int_{V}n_{\chi}$. After a characteristic time $\tau$ the system reaches annihilation equilibrium where $dN_{\chi}/dt\rightarrow 0$, which implies that $N_{\chi}\rightarrow\sqrt{C_{\rm cap}/C_{\rm ann}}$, leading to $\tau=1/\sqrt{C_{\rm cap}C_{\rm ann}}$.
For our forming protoplanet the $V_{\rm eff}\approx V_{\rm core}$, which typically has a $R\sim 3\,R_{\oplus}$, we thus have
$\tau\approx\sqrt{\frac{m_{\chi}\,10^{-30}\text{ cm}^{3}/\text{s}}{\langle\sigma_{\rm ann}v\rangle\text{ GeV}}}\text{ Myr}\,,$ (16)
which for DM masses around a GeV, and typical thermal annihilation cross sections is below the Myr time-scale. We therefore assume that the protoplanet reaches annihilation equilibrium, and the total rate of captured DM particles is directly converted to the DM luminosity, as discussed above.
Figure 2: Constraints on spin-dependent parameter space. The filled purple contours demonstrate the cross sections for which the formation of Jupiter stalls at $10\rm\,M_{\oplus}$, before gas is captured (see text for motivation of this benchmark). The green contours demonstrate the cross sections for which the formation of Jovian planets near the GC stalls at this same mass. Continuous contours assume that the environment in which the planet forms is $115\rm K$; dashed contours assume it is $20\rm K$. Reflection of light DM is considered for all but the outer purple $T=115\rm K$ contour. As we discussed above reflection is a model dependent phenomenon, and can be strongly suppressed in models with a long range force mediator. Finally, in the region below the orange dotted line DM evaporates from the forming planet as discussed in appendix A, given short-range interactions only. We also show constraints from direct detection (filled gray contours) and from the CMB Boddy and Gluscevic (2018) and MW-satelites Nadler et al. (2019) (dashed lines).
Figure 3: Constraints on spin-independent DM interactions from the existence of Jupiter, and expected sensitivity from the non-existence of Jovians in the GC, as well as direct detection and cosmology dependent constraints – as in Fig.2.
Implications
We study the DM fraction needed to halt the formation process for Jupiter in the solar system, using $\rho_{\rm DM}=0.42\,\rm GeV\,cm^{-3}$ and $v_{\rm DM}=270\rm km\,s^{-1}$, and find that $f_{\rm cap}$ remains approximately constant for the period of formation that we are interested in.
In Figs. 2 and 3 we show constraints on the spin-dependent and spin-independent scattering cross sections respectively from the fact that the formation of Jupiter in our solar system has not stopped at a mass of $10\rm M_{\oplus}$ as purple contours. As the temperature increase needed to satisfy (Disrupting Jovian Formation) does grow as a function of time, we expect that formation of Jovian planets in the GC halts at $10\rm M_{\oplus}$ in the parameter range indicated by the contours, and proceeds as normal below these contours. We also project constraints based on the observation of Jovian planets in the GC (for which we have assumed $\rho_{\rm DM}=10^{3}\,\rm GeV\,cm^{-3}$ and $v_{\rm DM}=10\rm\,km\,s^{-1}$) in chartreuse (note that the premature ending is already in tension with inferences of Jovian planets in the GC from lensing events Sahu et al. (2006); Janczak et al. (2010); Koshimoto et al. (2021)). We show these constraints and projections for two different benchmarks for the protostellar disk temperature; the more conservative $115$ K and $20$ K, for planets forming further away from the star. We note that this is not the only parameter degeneracy; most notably, the unavailability of nearby gas and dust can also cause a protoplanet to stop accreting, and the migration of a planet away from the star can likewise change its accretion rate. Therefore, while one can interpret the observation of a large gas giant as a constraint on DM, the non-observation cannot directly be interpreted as positive evidence.
For comparison we show current direct detection constraints as gray contours. Here we show the maximal reach in the sensitivity to the direct nuclear recoil process, which are least model dependent. The reach is dominated by the CRESST results at lower masses for spin-independent Angloher et al. (2023), and spin-dependent Angloher et al. (2022) cross sections, while at higher masses the LUX Akerib et al. (2017) results dominate for spin-dependent interactions, while the XENON experiment dominates the spin-independent sensitivity Aprile et al. (2023). For large spin-independent cross sections, we also show the XQC constraint Erickcek et al. (2007); Mahdawi and Farrar (2018). Note that searches based on the Migdal effect Zhang et al. (2022) could potentially reach lower DM masses, however, the results have large theoretical uncertainties Cox et al. (2023), and are not displayed here.
The sensitivity ceiling of direct detection experiments at large cross sections for light DM from Ref. Cappiello (2023) and for heavy DM from Refs. Kavanagh (2018); Cappiello et al. (2021) is shown. Since at the large cross sections considered the scaling with atomic mass number is uncertain Digman et al. (2019), we do not rescale the results for spin-dependent interactions and a more detailed analysis to obtain a precise ceiling values is needed. Note that in contrast the sensitivity of our constraints and suggested search is only affected at cross sections above $10^{-22}\text{cm}^{2}$, where large drift times would require a more detailed analysis of the DM distribution Acevedo et al. (2021); Leane and Smirnov (2022) and annihilation outside the core.
We also show complementary cosmology-dependent constraints from the CMB Boddy and Gluscevic (2018). DM which scatters with a velocity-independent cross-section can also be constrained from the lack of MW satellites produced in such models Nadler et al. (2019). However, as the DM scattering rates could be different in the early universe Elor et al. (2023), those bounds are not as universal as the shown late-time sensitivity.
Lastly, as a result of DM heat injection, the proto-planet will respond by radiating away more energy. Even if the heating effect is not significant enough to halt the formation of the gas envelope, this effect leads to an increase in luminosity in high DM-density environments. The luminosity of protoplanet formation peaks in the near-infrared and can be determined by future observations. This allows a direct comparison of the observed peak luminosity, realised at about $2$ Myr, and that predicted by theoretical models.
Discussion
In this paper, we have studied the inhibition of the formation of Jupiter in our solar system and Jovian planets in the Milky Way galaxy by the capture and annihilation of DM. We have shown that the heat generated by DM annihilation can increase the temperature of the protoplanet above the temperature at which the gas mixture evaporates, and determined the corresponding constraints on the parameter space of spin-(in)dependent DM. These constraints are competitive for relatively strongly interacting DM candidates, above the “ceiling” of direct detection constraints. We have shown that these constraints depend on the ambient temperature at formation, with stronger effects for planets forming further away from the star.
We have also estimated the projected sensitivities from the existence of Jovian planets in the GC, where the DM density is higher and the typical velocity dispersion lower. Future expolanet observations by JWST and Roman may be used to test these scenarios. Moreover, as we explain in section Implications, for a particular DM (sub-component) mass and cross section the formation of Jovian planets is affected as a function of the distance from the GC, enabling a differential detection of the heating signal.
In this work, we have estimated the sensitivity of protoplanets to DM injections using a semi-analytical calculation, with input from simulations in the absence of these injections. Future work should focus on the embedding of DM heat in planetary formation simulations, such that more realistic projections can be obtained. Here we have focused on the core-accretion gas-capture theory of giant planet formation. This is the simplest scenario for the formation of Jovian planets. Alternative scenarios include dynamical processes and migration within the stellar system. Since we focus on the final stages of formation, we expect such alternative scenarios to be qualitatively similar.
Acknowledgements
We thank Chris Cappiello, Malcolm Fairbairn, Scott Gaudi, Rebecca Leane, Dave McKeen, and Nirmal Raj for useful conversations and feedback on the manuscript. DC is supported by the STFC under Grant No. ST/T001011/1.
Appendix A Dark matter evaporation
The condition (8) relies on the fact that hydrogen evaporates for temperatures above $T_{\rm max}$. Therefore, one would naively say that DM candidates with masses below the mass of hydrogen would evaporate at these temperatures too. However, unlike hydrogen, as we demonstrate below the DM particles are concentrated in the core of the objects. Thus, in the short mean free path regime the lower number density of DM particles in the envelope implies a lower flux of escaping particles (Disrupting Jovian Formation). Here we estimate the evaporation rate of DM particles of a given mass, and compare it to the capture rate.
The DM number density within the protoplanet is obtained by integrating the radial equation Eq. (11) form Ref. Leane and Smirnov (2022), under the assumption that the envelope has constant density $\rho_{E}$ (at $t=\rm Myr$, $\sim 3\times 10^{-8}\rm g\,cm^{-3}$) and the core has a constant density $\rho_{\rm core}$ (at $t=\rm Myr$, $\sim 2\rm g\,cm^{-3}$). We distinguish two regimes, the isothermal regime, where the DM mean free path $\lambda$ in the object is of the order of the object, and DM thermalises with the core at a constant temperature. The second regime where local thermal equilibrium (LTE) is applicable when the DM mean free path is smaller than the size of the object. Thus we have the total DM profile:
$n_{\chi}=\begin{cases}g(r)/N_{0}\,\exp\left(-F(r)\right)\quad\text{if }\lambda<R_{E}\\ 1/N_{0}\exp(-m_{\chi}\Phi(r)/T(r))\quad\text{if }\lambda>R_{E}\end{cases}\,,$ (17)
where the radius is normalised to the envelope size $r=R/R_{E}$, $\Phi(r)$ is the gravitational potential, $T(r)$ the temperature profile, the normalisation factor $N_{0}$ ensured that we have $N_{\chi}=4\pi\int_{0}^{R_{E}}r^{2}n_{\chi}(r)$, and the LTE functions are given by
$\frac{F(r)}{\frac{4}{3}\pi G_{N}m_{\chi}R_{E}^{2}}=\frac{r^{2}\rho_{\rm core}}{T_{\rm core}}\theta\left(\frac{R_{\rm core}}{R_{E}}-r\right)+\theta\left(r-\frac{R_{\rm core}}{R_{E}}\right)$
×$\times\frac{\left(1/2\,r^{2}\rho_{E}+\left(\frac{R_{\rm core}}{R_{E}}\right)^{3}(\rho_{\rm core}-\rho_{E})(\frac{R_{E}}{R_{\rm core}}-\frac{1}{r})\right)}{T_{E}}\,,$ (18)
$g(r)=\theta\left(r-\frac{R_{\rm core}}{R_{E}}\right)\left(\frac{T_{E}}{T_{\rm core}}\right){}^{\frac{1}{2}\left(\frac{m_{\chi}}{m_{\text{p}}}+1\right)^{-3/2}-1}$
$+\theta\left(\frac{R_{\rm core}}{R_{E}}-r\right)\,,$ (19)
where $\theta$ indicates the Heaviside function. For the regime where the mean free path of DM becomes comparable to SM particles the above equation provides us with the DM density within the last scattering surface, to which we apply the Jeans’ escape criterion, demanding that the total surface evaporation rate is below the capture rate. While for the isothermal, and transition regime we use the formalism of Ref. Gould (1987, 1990) to determine the evaporation mass comparing the evaporation rate and the time-scale needed to reach annihilation equilibrium assuming a core temperature of $3000$ K. This leads to the evaporation contour, shown by the dotted lines in Figs. 23. Note that this line is only marking to lowest mass of the DM retained in the object if DM interacts through contact interactions only. In the presence of attractive long range forces, the evaporation contours are significantly moved to lower DM masses, as discussed in Ref. Acevedo et al. (2023).

### v3

@@
 The existence of Jupiter in our solar system can therefore be used to infer constraints on dark matter with relatively strong interaction cross sections.
-We derive novel constraints on the cross section for both spin-dependent and spin-independent dark matter.
+In the case of spin-dependent dark matter, we derive novel constraints beyond the reach of current direct detection experiments.
 We highlight the possibility of a positive detection using future observations by JWST, which could reveal strongly varying planet morpholoiges close to our Galactic Center.
@@
 Our technical capabilities to detect planets outside our solar system have developed dramatically.
-From the first indirect evidence for an exoplanet Mayor and Queloz (1995), to direct imaging Chauvin et al.
-(2005) and precision atmospheric observations Martins et al.
-(2015), we have now 5.5× $5.5\times\,10^{3}$ detected exoplanets, and $\mathcal{O}(10^{4})$ candidates awaiting confirmation cat (2020).
-Especially with the onset of high performance infra-red astronomy and the JWST Brande et al.
-(2019), and Roman Johnson et al.
-(2020) space missions low temperature targets, such as exoplanets can be studied across the galaxy, all the way to the Galactic Center (GC).
+From the first indirect evidence for an exoplanet 1995Natur.378..355M, to direct imaging 2005A&A...438L..29C and precision atmospheric observations 2015A&A...576A.134M, we have now 5.5× $5.5\times\,10^{3}$ detected exoplanets, and $\mathcal{O}(10^{4})$ candidates awaiting confirmation catalog.
+With the onset of high performance infra-red astronomy and the JWST Brande_2019, and Roman Johnson_2020 space missions low temperature targets, such as exoplanets can be studied across the galaxy, all the way to the Galactic Center (GC).
 This development opens up a new window to our galaxy, and allows us to study exoplanets in various astrophysical environments.
-Dark Matter (DM) as the dominant mass source in our galaxy is well mapped by stellar kinematics, and despite uncertainties displays a density profile that raises towards the GC Benito et al.
-(2021).
-In the search for this ellusive substance a number of studies have considered DM capture in celestial object ranging from the Earth and the Sun Batell et al.
-(2010); Pospelov et al.
-(2008); Pospelov and Ritz (2009); Rothstein et al.
-(2009); Chen et al.
-(2009); Schuster et al.
-(2010a, b); Bell and Petraki (2011); Feng et al.
-(2016a); Kouvaris and Tinyakov (2010a); Feng et al.
-(2016b); Allahverdi et al.
-(2017); Leane et al.
-(2017); Arina et al.
-(2017); Albert et al.
-(2018a, b); Nisa et al.
-(2019); Niblaeus et al.
-(2019); Cuoco et al.
-(2020); Serini et al.
-(2020); Acevedo et al.
-(2021); Mazziotta et al.
-(2020); Bell et al.
-(2021), Jupiter Batell et al.
-(2010); Leane and Linden (2021); Li and Fan (2022); French and Sher (2022); Ray (2023), Brown Dwarfs Leane and Smirnov (2021); Leane et al.
-(2021), Uranus Mitra (2004), White Dwarfs Garani et al.
-(2023), Neutron Stars Goldman and Nussinov (1989); Gould et al.
-(1990); Kouvaris (2008); Bertone and Fairbairn (2008); de Lavallaz and Fairbairn (2010); Kouvaris and Tinyakov (2010b); McDermott et al.
-(2012); Kouvaris and Tinyakov (2011a); Guver et al.
-(2014); Bramante et al.
-(2013); Bell et al.
-(2013); Bramante et al.
-(2014); Bertoni et al.
-(2013); Kouvaris and Tinyakov (2011b); McCullough and Fairbairn (2010); Angeles Perez-Garcia and Silk (2015); Bramante (2015); Graham et al.
-(2015); Cermeno et al.
-(2016); Graham et al.
-(2018); Acevedo and Bramante (2019); Janish et al.
-(2019); Krall and Reece (2018); Baryakhtar et al.
-(2017); Raj et al.
-(2018); Bell et al.
-(2018); Chen and Lin (2018); Dasgupta et al.
-(2019); Hamaguchi et al.
-(2019); Camargo et al.
-(2019); Bell et al.
-(2019); Acevedo et al.
-(2020); Joglekar et al.
-(2019, 2020); Bell et al.
-(2020); Garani et al.
-(2020); Leane et al.
-(2021); Bramante et al.
-(2022a); Coffey et al.
-(2022), and other stars Freese et al.
-(2009); Taoso et al.
-(2008); Ilie et al.
-(2020a, b); Lopes and Lopes (2021); Ellis (2021); Croon and Sakstein (2023), as well as recently Exoplanets Leane and Smirnov (2021); Bramante et al.
-(2022b) (for a recent review, see Bramante and Raj (2023)).
+Dark Matter (DM) as the dominant mass source in our galaxy is well mapped by stellar kinematics, and despite uncertainties displays a density profile that raises towards the GC Benito:2020lgu.
+In the search for this ellusive substance a number of studies have considered DM capture in celestial object ranging from solar system objects such as the Earth, Sun Batell:2009zp; Pospelov:2007mp; Pospelov:2008jd; Rothstein:2009pm; Chen:2009ab; Schuster:2009au; Schuster:2009fc; Bell_2011; Feng:2015hja; Kouvaris:2010; Feng:2016ijc; Allahverdi:2016fvl; Leane:2017vag; Arina:2017sng; Albert:2018jwh; Albert:2018vcq; Nisa:2019mpb; Niblaeus:2019gjk; Cuoco:2019mlb; Serini:2020yhb; Acevedo:2020gro; Mazziotta:2020foa; Bell:2021pyy, Jupiter Batell:2009zp; Leane:2021tjj; Li:2022wix; French:2022ccb; Ray:2023auh, and Uranus Mitra:2004fh, to Brown Dwarfs Leane:2020wob; Leane:2021ihh, White Dwarfs Garani:2023esk, Neutron Stars Goldman:1989nd; Gould:1989gw; Kouvaris:2007ay; Bertone:2007ae; deLavallaz:2010wp; Kouvaris:2010vv; McDermott:2011jp; Kouvaris:2011fi; Guver:2012ba; Bramante:2013hn; Bell:2013xk; Bramante:2013nma; Bertoni:2013bsa; Kouvaris:2010jy; McCullough:2010ai; Perez-Garcia:2014dra; Bramante:2015cua; Graham:2015apa; Cermeno:2016olb; Graham:2018efk; Acevedo:2019gre; Janish:2019nkk; Krall:2017xij; Baryakhtar:2017dbj; Raj:2017wrv; Bell:2018pkk; Chen:2018ohx; Dasgupta:2019juq; Hamaguchi:2019oev; Camargo:2019wou; Bell:2019pyc; Acevedo:2019agu; Joglekar:2019vzy; Joglekar:2020liw; Bell:2020jou; Garani:2020wge; Leane:2021ihh; Bramante:2021dyx; Coffey:2022eav; Nguyen:2022zwb; Alvarez:2023fjj, and other stars Freese:2008hb; Taoso:2008kw; Ilie:2020iup; Ilie:2020nzp; Lopes:2021jcy; Ellis:2021ztw; DCandJS, as well as recently Exoplanets Leane:2020wob; Bramante:2022pmn.
+For a recent review of searches in compact stars, see Bramante:2023djs.
 Many of the above studies rely on an energy injection from DM particles that annihilate after being captured by the celestial object.
@@
 Jovian exoplanets have been identified as ideal candidates for DM capture in Ref.
-Leane and Smirnov (2021).
+Leane:2020wob.
 The advantage is their relatively large surface gravity, combined with rather low core temperatures, leading to DM with low masses to be retained in the object.
 Their hydrogen-rich composition makes them kinematically well suited for capturing DM particles around the GeV scale, and especially particles with spin-dependent nuclear interactions.
-The differential measurement of Jovian planet and Brown Dwarf temperatures close to the GC has been identified as a promising tool to detect heating from annihilating DM  Leane and Smirnov (2021).
+The differential measurement of Jovian planet and Brown Dwarf temperatures close to the GC has been identified as a promising tool to detect heating from annihilating DM  Leane:2020wob.
 In this work we propose a more radical possibility.
@@
 Thus, we would expect to have varying properties of Jovian planets across the galaxy, or their total absence under certain circumstances.
-Figure 1: Benchmark for spin-dependent dark matter: $m_{\rm DM}=10$ GeV, $\sigma_{p}=4\times 10^{-28}\,\rm cm^{2}$.
-The red line shows the dark matter capture fraction at the time when the planet is $10\,\rm M_{\oplus}$ and the black continuous and dashed lines show the fractions needed to halt formation through gas capture.
-This fraction depends on the temperature of the surrounding gas, and thus on the proximity to the host star.
-In Fig.
-1 we show the capture efficiency as a function of distance from the GC assuming an NFW profile and a Maxwellian velocity dispersion.
-As shown here, at a given distance from the host star (corresponding to a given external temperature $T_{\rm ext}$), planet formation can be halted in the GC while it does occur some distance away from it.
-This would imply that towards the GC, Jovian planets appear to occur further closer to the host stars, modulo migration after formation.
-This observation provides a new and radically different search for a DM heating signal within our galaxy.
-This distinct phenomenology can be targeted by future observations of exoplanets leveraging the advances in the development of infrared telescopes.
-Finally, the mere existence of Jupiter in our solar system allows us to set new constraints on the parameter space of DM.
+Figure 1: Captured DM fraction in a benchmark scenario for spin-dependent dark matter: $m_{\rm DM}=5$ GeV, $\sigma_{p}=10^{-27}\,\rm cm^{2}$.
+Dark matter density and velocity as in the solar system.
+The red line gives the fraction of the dark matter flux through the planet that is captured by the core, the cyan line gives the fraction captured by the envelope, and the dashed line the total captured fraction.
+Above the black line labeled $T>T_{\rm max}$ dark matter annihilation heats the planet above the temperature defined by Eq.
+(2), which halts the formation process.
+The mere existence of Jupiter in our solar system allows us to set new constraints on the parameter space of DM, which are especially competitive in the case of spin-dependent interactions.
 We find that future observations of exoplanets near the GC, enabled by the JWST and future infrared telescopes will further test yet unconstrained parameter space of DM candidates across a wide range of masses and interaction cross sections.
 Planet formation
-We focus on giant, Jovian planets, and the formation process called core-accretion gas-capture Hubickyj et al.
-(2005), which proceeds in two stages.
+We focus on giant, Jovian planets, and the formation process called core-accretion gas-capture 2005Icar..179..415H, which proceeds in two stages.
 First, a solid core made of heavier elements, such as ice and silica forms within a fraction of a Myr.
@@
 The proto-planet then accumulates a gas envelope at an increasing rate, reflecting its increasing mass.
-The gas can quickly fill up the Bondi sphere of the planet, which subsequently needs to cool via the Kelvin-Helmholtz mechanism in order to accrete further Pollack et al.
-(1996); Adams and Batygin (2022), until eventually the protoplanet reaches a threshold mass at which runaway accretion takes place.
-Therefore, during the gas accretion period before runaway (which takes place on the scale of a Myr  D’Angelo et al.
-(2021)) the system remains in a quasi-equilibirum state where gas accretion can only proceed once the excess heat that is released from the gravitational binding energy has been radiated away.
+The gas can quickly fill up the Bondi sphere of the planet, which subsequently needs to cool via the Kelvin-Helmholtz mechanism in order to accrete further 2022DPS....5410202A.
+Therefore, during the gas accretion period (which takes place on the scale of a Myr  2021Icar..35514087D) the system remains in a quasi-equilibirum state where gas accretion can only proceed once the excess heat that is released from the gravitational binding energy has been radiated away.
 It is then not surprising that an additional heat source can disrupt this fragile equilibrium and lead to drastically different outcomes.
@@
 $\begin{split}\dot{Q}-L_{\rm DM}(R)&=-\frac{dU}{dt}\\ 4\pi R^{2}\times\sigma_{\rm SB}T^{4}-L_{\rm DM}(R)&=c_{1}\frac{GM^{2}}{R^{2}}\frac{dR}{dt}\end{split}$ (1)
-where $\dot{Q}$ is the radiative heat loss, $dU/dt$ the change in gravitational potential energy, $L_{\rm DM}(R)$ is the $R$ dependent heat injection due to DM annihilation, which we call DM luminosity (note that this is not necessarily directly observable as luminosity of the protoplanet), and $1/2\leq c_{1}\leq\infty$ is a factor which depends on the internal mass distribution (for a uniform sphere, $c_{1}=3/5$).
+where $\dot{Q}$ is the radiative heat loss, $dU/dt$ the change in gravitational potential energy, $L_{\rm DM}(R)$ is the $R$ dependent heat injection due to DM annihilation, which we call DM luminosity, and $1/2\leq c_{1}\leq\infty$ is a factor which depends on the internal mass distribution (for a uniform sphere, $c_{1}=3/5$).
 Initially, dark matter heat injections will raise the temperature of the protoplanet, increasing its radiative heat loss.
-However, above a certain temperature the mixture of hydrogen and helium gas acquires enough thermal energy such that particles near its surface reach the escape energy and become unbound.
-This process will be in competition with the mass accretion rate, which we take from numerical simulations Lissauer et al.
-(2009).
-To quantify the particle loss rate we consider the generalisation of the Jeans escape rate $\Gamma^{0}_{\rm J}(T)$ Jeans (2009) by including the effect of a chemical potential difference
-$\Gamma_{\rm J}(T)=\Gamma^{0}_{\rm J}(T)\exp\left(-\frac{\Delta\mu}{N\,T}\right)$
-$=\frac{nv}{2\sqrt{\pi}}\left(1+\frac{v_{\rm esc}^{2}}{v^{2}}\right)\exp\left(-\frac{E_{\rm esc}}{T}\right),$ (2)
-where $n$ is the number density of gas particles, $v_{\rm esc}$ the surface escape velocity, $v=\sqrt{2T/m}$, and the escape energy that takes into account the difference in chemical potentials
-$E_{\rm esc}=\frac{G_{N}Mm}{r}+\frac{\Delta\mu}{N}=\frac{m\,v_{\rm esc}^{2}}{2}+\frac{\Delta\mu}{N}\,.$ (3)
-Here, $G_{N}$ is the Newton constant, $M$ the mass of the forming planet, $m$ the mass of the evaporating particle, and $\Delta\mu/N$ is the difference in chemical potentials per particle.
-It is intriguing that the chemical potential is relevant in this case, reflecting the special situation of evaporation of particles from the planetary envelope into the surrounding gas of the protoplaneary disk, which has a comparable particle number density and temperature.
-Given that $\mu_{1}$ is the chemical potential in the envelope and $\mu_{2}$ that of the protoplaneraty disk, the difference in the chemical potentials $\Delta\mu$ is given by
-$\mu_{1}-\mu_{2}=U_{1}-U_{1}+P_{1}V_{1}-P_{2}V_{2}+T_{2}S_{2}-T_{1}S_{1}$ (4)
-where we assumed that the gravitational energies of the particles are of the same order.
-We further assume that the particles behave as ideal gas, and thus the internal energy is $U_{i}=C_{V}N_{i}T_{i}$, $P_{i}V_{i}=N_{i}T_{i}$, and the entropy is given by the Sackur-Tetrode formula $S_{i}=N_{i}\kappa_{i}$, with
-$\kappa_{i}=\left(\frac{5}{2}+\log\left[\frac{V_{i}}{N_{i}}\left(\frac{4\pi mU_{i}}{3N_{i}}\right)^{3/2}\right]\right)\,.$ (5)
-Furthermore, we assume that the gases are monoatomic, and hence the heat capacity $C_{V}\approx 3/2$.
-We can therefore express the chemical potential difference per envelope particle as
-$\frac{\Delta\mu}{N_{1}}=C_{V}\left(T_{1}-\frac{N_{2}}{N_{1}}T_{2}\right)+T_{1}-\frac{N_{2}}{N_{1}}T_{2}+\bar{\kappa}\left(T_{2}\frac{N_{2}}{N_{1}}-T_{1}\right)$
-$\approx\left(1+C_{V}-\bar{\kappa}\right)\left(T_{1}-T_{2}\right)=\frac{3}{2}\log\left[\frac{4\pi m\bar{u}}{3\bar{n}^{2/3}}\right]\left(T_{2}-T_{1}\right)\,.$ (6)
-Where we have used the assumption that $N_{1}\approx N_{2}$, and as the logarithm is a slowly varying function, the argument can be replaced by the average value of the internal energy per particle $\bar{u}=\bar{U}/\bar{N}\approx C_{V}\bar{T}$.
-Thus, we observe that in the case that the surrounding gas is hotter than the envelope an additional suppression slows particle evaporation down.
-Now we compare the escape rate with the rate of growth of the planet in the absence of dark matter heat $\dot{M}$, which we take from simulation results presented in Lissauer et al.
-(2009).
-Equating those rates
-$A_{\rm planet}\,\Gamma_{\rm J}(T)|_{T=T_{\rm max}}=\mathcal{O}(1)\,\dot{M}$ (7)
-where $A_{\rm planet}$ is the surface from which evaporation takes place, we find that the result is insensitive to the value of the numerical $\mathcal{O}(1)$ factor in the equation, since it can not compete with the exponential in the escape rate (see also D’Angelo et al.
-(2021), which is largely consistent but presented slower accretion, implying smaller DM injections are needed to match it).
-For an ambient temperature of $115$ K of the surrounding gas Lissauer et al.
-(2009), this treatment gives us hydrogen escape temperatures of $T_{\rm max}\sim 85$ K for the envelope.
-For the core the assumption above has to be modified given that the particle number density in the core is much higher than in the surrounding gas.
-We get $T_{\rm max}\sim 10^{3}$ K for the core, a result that is not significantly affected by the ambient gas properties for the systems we study.
-Thus, the maximum heat loss happens just below $T_{\rm max}$ and we estimate that DM heat injection halts further gas capture if
-$\begin{split}L_{\rm DM}&\geq 4\pi R^{2}\sigma T_{\rm max}^{4}\\ &\geq 4\times 10^{-4}\times\left(\frac{R}{10^{3}R_{p}}\right)^{2}\,\left(\frac{T_{\rm max}}{80\rm K}\right)^{4}\rm\,L_{\odot}\end{split}$ (8)
+However, above a certain temperature the mixture of hydrogen and helium gas acquires enough thermal energy such that particles near its surface reach the escape velocity.
+We estimate that this happens for temperatures above which
+$\Gamma_{\rm J}(T)=\left.\frac{nv}{2\sqrt{\pi}}\left(1+\frac{v_{\rm esc}^{2}}{v^{2}}\right)\exp\left(-\frac{v_{\rm esc}^{2}}{v^{2}}\right)\right|_{T=T_{\rm max}}=C\dot{M}$ (2)
+where $\Gamma_{\rm J}$ is the flux of escaping particles given by the Jeans escape formula jeans_2009, with $n$ the number density of gas particles, $v_{\rm esc}$ the surface escape velocity, $v=\sqrt{2T/m}$, with $m$ being the mass of the gas particle, and $\dot{M}$ is the rate of growth of the planet in the absence of dark matter heat, which we take from simulation results presented in Lissauer:2008hn; d2021growth.
+Equation (2) overestimates the actual loss rate because it assumes the tail of the velocity distribution is replenished instantaneously, Therefore, to be conservative, we take $C=10$.
+This gives us maximum hydrogen escape temperatures of $T_{\rm max}\sim 8$ K and $10^{3}$ K in the envelope and in the core respectively.
+Thus, the maximum heat loss happens just below $T_{\rm max}$ and we estimate that DM heat injection halts collapse if
+$\begin{split}L_{\rm DM}&\geq 4\pi R^{2}\sigma T_{\rm max}^{4}\\ &\geq 4\times 10^{-8}\times\left(\frac{R}{10^{3}R_{p}}\right)^{2}\,\left(\frac{T_{\rm max}}{8\rm K}\right)^{4}\rm\,L_{\odot}\end{split}$ (3)
 where $R_{p}=7\times 10^{5}$ km is the radius of Jupiter (the protoplenatary envelope has radius $\sim 10^{3}R_{p}$), and where $L_{\odot}=2\times 10^{36}\,\rm GeV\,s^{-1}$ is the luminosity of the Sun.
 To investigate whether Eq.
-(8) is satisfied in the evolution of the planet, we use the simulation results presented in D’Angelo et al.
-(2021).
+(3) is satisfied in the evolution of the planet, we use the simulation results presented in d2021growth.
 We will now examine how the DM luminosity depends on the properties of the capturing object, and the DM particle model.
@@
 We thus can write
-$L_{\rm DM}=m_{\rm DM}C_{\rm cap}=m_{\rm DM}f_{\rm cap}\Phi\pi R^{2}$ (9)
-where $C_{\rm cap}$ is the capture rate, and $f_{\rm cap}$ the captured fraction of the DM flux $\Phi$ passing through the object, given by
-$\Phi=\,v_{\rm DM}\sqrt{\frac{8}{3\pi}}\frac{\rho_{\rm DM}}{m_{\rm DM}}\left(1+\frac{3}{2}\frac{v_{\rm esc}^{2}}{v_{\rm DM}^{2}}\right)\,.$ (10)
+$L_{\rm DM}=m_{\chi}C_{\rm cap}=m_{\chi}f_{\rm cap}\Phi$ (4)
+where $m_{\chi}$ is the DM mass, $C_{\rm cap}$ is the capture rate, and $f_{\rm cap}$ the captured fraction of the DM flux $\Phi$ passing through the object, given by
+$\Phi=v_{\rm DM}\sqrt{\frac{8}{3\pi}}\frac{\rho_{\rm DM}}{m_{\chi}}\left(1+\frac{3}{2}\frac{v_{\rm esc}^{2}}{v_{\rm DM}^{2}}\right)\,.$ (5)
 Then, we have
-$\begin{split}L_{\rm DM}\sim&7\times 10^{-4}f_{\rm cap}\left(\frac{R}{10^{3}R_{p}}\right)^{2}\\ &\times\left(\frac{v_{\rm DM}}{270\rm km\,s^{-1}}\right)\left(\frac{\rho_{\rm DM}}{0.42\,\rm GeV\,cm^{-3}}\right)\,L_{\odot}\end{split}$ (11)
-which only depends on the dark matter mass through $f_{\rm cap}$. It is seen that this may indeed exceed (8).
+$\begin{split}L_{\rm DM}\sim&2\times 10^{-9}f_{\rm cap}\left(\frac{R}{10^{3}R_{p}}\right)\left(\frac{M}{10M_{\oplus}}\right)\\ &\times\left(\frac{v_{\rm DM}}{270\rm km\,s^{-1}}\right)^{-1}\left(\frac{\rho_{\rm DM}}{0.42\,\rm GeV\,cm^{-3}}\right)\,L_{\odot}\end{split}$ (6)
+independently of the dark matter mass. It is seen that this may exceed (3) for greater $M$ or $\rho_{\rm DM}$.
 Dark matter capture
 Given a certain DM mass and velocity, the DM particle needs a certain number of scatterings to drop below the escape velocity of a given object.
-This will in general also depend on the efficiency of the energy loss, determined by the mass ratio of the DM particle and the target nucleus – DM is not sufficiently slowed down when this mass ratio is too large.
-Furthermore, in the case of DM particles with masses similar to or below the target nucleus mass, DM particles might be reflected off the object before capture can occur, as discussed in detail in Ref.
-Leane and Smirnov (2023a, b).
-To compute the fraction of the DM particles that are captured in a given object we use the Asteria Leane and Smirnov (2023b, a) package, which takes all the above mentioned phenomena into account.
+This will in general also depend on the efficiency of the energy loss, determined by the mass ratio of the DM particle and the target nucleus.
+Furthermore, in the case of DM particles with masses similar of below the target mass, DM particles might be reflected off the object before capture can occur, as discussed in detail in Ref.
+RKLandJS; asteria.
+To compute the fraction of the DM particles that are captured in a given object we use the Asteria asteria; RKLandJS package, which takes all the above mentioned phenomena into account.
 In this work we focus on DM interaction with nuclei and distinguish two broad DM model classes.
@@
 where $\mu(m_{\chi},m_{\rm SM})$ is the reduced mass of a DM and a SM particle of mass $m_{\rm SM}$.
-Note that it has been pointed out that at cross sections $\sigma_{\chi N}>10^{-26}\text{cm}^{2}$ the scattering is not described by interactions of point like particles, and the coherent scattering picture breaks down Digman et al.
-(2019); Cappiello et al.
-(2021).
+Note that it has been pointed out that at cross sections $\sigma_{\chi N}>10^{-26}\text{cm}^{2}$ the scattering is not described by interactions of point like particles, and the coherent scattering picture breaks down Digman:2019wdm; Cappiello:2020lbk.
 Rather, at cross sections larger than this threshold the interaction can be assumed as a total geometric scattering cross section with the entire nucleus without an implied scaling with atomic number.
@@
 For example in the case of $a_{p}=1$, and $a_{n}=0$, the hydrogen scattering cross section is equal to the DM scattering cross section on protons $\sigma_{SD}=\sigma_{p}$, which we will use as a benchmark model.
-To compute the capture of DM in the protoplanet we introduce a simplifying effective framework and use the Asteria Leane and Smirnov (2023b, a) package to compute capture in two regimes depending on where in the object we expect the DM to be captured.
-In the first regime the DM capture takes place on the scale of the entire object and the average chemical composition of the protoplanet is taken as an input for Asteria, while in the other regime the capture takes place in the envelope.
-In the second case only the chemical composition of the envelope is used, but the total mass of the protoplanet determines the escape velocity.
-To distinguish the regimes we define the effective capture length $\lambda_{\rm cap}$ and compare it to the size of the protoplanet, which is dominated by the size of the envelope.
-$\lambda_{\rm cap}=\lambda_{\rm mfp}\,N_{\rm cap}\text{ with }\lambda_{\rm mfp}=\left(\sigma_{\rm SI/SD}\,n_{\rm SM}\right)^{-1}\,,$ (14)
-$N_{\rm cap}=\frac{\log\left(v_{\text{esc}}^{2}/(v_{\chi}^{2}+v_{\text{esc}}^{2})\right)}{\log\left(1-2\mu(1-\mu)^{-2}\right)}\,,$
-where $N_{\rm cap}$ is the number of scatterings required to capture a DM particle Mack et al.
-(2007), $v_{\text{esc}}$ the protoplanets’ escape velocity, $v_{\chi}$ the DM impact velocity, and $\mu=m_{\rm DM}/m_{\rm SM}$ is the mass ratio between the DM and the average SM particle in the protoplanet.
-Note that this estimate provides an upper bound on the capture length of DM as a straight-line trajectory is assumed.
-To determine the chemical composition we model the protoplanet in a simplified manner assuming that it consists of an icy core are an outer gas envelope, both are assumed to have a homogeneous average density.
+We model the protoplanet in a simplified manner assuming that it consists of an icy core are an outer gas envelope, both are assumed to have a homogeneous average density.
 We use the mass and radius from Ref.
-Lissauer et al.
-(2009) as time dependent input quantities.
+Lissauer:2008hn as time dependent input quantities.
 The outer envelope has a large radius and a low density.
-We model the envelope to consist of 75% hydrogen and 25% helium.
-The core is assumed to have a $58/42$ ice ($\rm H_{2}O$) to silica ($\rm SiO_{2}$) mass ratio D’Angelo et al.
-(2021), which makes Oxygen the most abundant element.
-This implies that in the regime where the core contributes to the capture, optimal capture happens for slightly heavier DM candidates, due to more efficient energy transfer between the SM and DM particles of similar mass.
+We model the envelope to consist of 75% hydrogen and 25% helium and compute the fraction of dark matter that is captured.
+We find that only at substantial cross sections above $10^{-28}\text{cm}^{2}$ a substantial fraction of the DM particles are captured.
+The most efficient capture rate takes place in the GeV mass range, when the DM mass matches the hydrogen target mass.
+The core of the protoplanet is much more dense, which has the potential to increase the DM capture rate.
+We model the core as having a $58/42$ ice ($\rm H_{2}O$) to silica ($\rm SiO_{2}$) mass ratio 2021Icar..35514087D, which makes Oxygen the most abundant element.
+This implies that optimal capture happens for slightly heavier DM candidates, due to more efficient energy transfer between the SM and DM particles of similar mass.
 For the spin-dependent calculation, we need to estimate the mass fraction which is relevant for spin-dependent scattering. Thus we take the isotope fractions and properties of the elements in Tab. 1.
@@
 $\langle S_{n}\rangle$ 0.508 0.204 0
-Lastly, we note that light DM candidates can be reflected, such that the capture efficiency remains below one even at large cross sections Leane and Smirnov (2023a).
-On the other hand, in models with a long range force mediator, the escape velocity in the medium can be increased, which in turn suppressed the reflection probability Acevedo et al.
-(2023).
-We discuss the implications for the cross section sensitivity estimate in the following section.
+Lastly, we note that light DM candidates can be reflected, such that the capture efficiency remains below one even at large cross sections RKLandJS.
+As we will see below, this has implications for the sensitivity to subfractions of DM.
 Dark matter annihilation
@@
 Figure 2: Constraints on spin-dependent parameter space.
-The filled purple contours demonstrate the cross sections for which the formation of Jupiter stalls at $10\rm\,M_{\oplus}$, before gas is captured (see text for motivation of this benchmark).
-The green contours demonstrate the cross sections for which the formation of Jovian planets near the GC stalls at this same mass.
-Continuous contours assume that the environment in which the planet forms is $115\rm K$; dashed contours assume it is $20\rm K$.
-Reflection of light DM is considered for all but the outer purple $T=115\rm K$ contour.
-As we discussed above reflection is a model dependent phenomenon, and can be strongly suppressed in models with a long range force mediator.
+The filled contour demonstrates the cross sections for which the formation of Jupiter stalls at $30\rm\,M_{\oplus}$.
+The dashed contours demonstrate the cross sections for which the formation of Jovian planets near the GC stalls at $10\rm\,M_{\oplus}$ (see text for motivation of these conditions).
+The reflection of light DM candidates implies that the capture efficiency asymptotes.
+Hence, for a subfraction $f_{\rm DM}=10^{-3}$ (shown in chartreuse) there are no constraints below $m_{\chi}\sim 100$ MeV (the asymptotic value is below the value needed to realise (3)).
 Finally, in the region below the orange dotted line DM evaporates from the forming planet as discussed in appendix A, given short-range interactions only.
-We also show constraints from direct detection (filled gray contours) and from the CMB Boddy and Gluscevic (2018) and MW-satelites Nadler et al.
-(2019) (dashed lines).
-Figure 3: Constraints on spin-independent DM interactions from the existence of Jupiter, and expected sensitivity from the non-existence of Jovians in the GC, as well as direct detection and cosmology dependent constraints – as in Fig.2.
+Figure 3: Constraints on spin-independent DM interactions from the existence of Jupiter, and expected sensitivity from the non-existence of Jovians in the GC, as well as current direct detection constraints– as in Fig.2.
 Implications
-We study the DM fraction needed to halt the formation process for Jupiter in the solar system, using $\rho_{\rm DM}=0.42\,\rm GeV\,cm^{-3}$ and $v_{\rm DM}=270\rm km\,s^{-1}$, and find that $f_{\rm cap}$ remains approximately constant for the period of formation that we are interested in.
-In Figs.
-2 and 3 we show constraints on the spin-dependent and spin-independent scattering cross sections respectively from the fact that the formation of Jupiter in our solar system has not stopped at a mass of $10\rm M_{\oplus}$ as purple contours.
-As the temperature increase needed to satisfy (Disrupting Jovian Formation) does grow as a function of time, we expect that formation of Jovian planets in the GC halts at $10\rm M_{\oplus}$ in the parameter range indicated by the contours, and proceeds as normal below these contours.
-We also project constraints based on the observation of Jovian planets in the GC (for which we have assumed $\rho_{\rm DM}=10^{3}\,\rm GeV\,cm^{-3}$ and $v_{\rm DM}=10\rm\,km\,s^{-1}$) in chartreuse (note that the premature ending is already in tension with inferences of Jovian planets in the GC from lensing events Sahu et al.
-(2006); Janczak et al.
-(2010); Koshimoto et al.
-(2021)).
-We show these constraints and projections for two different benchmarks for the protostellar disk temperature; the more conservative $115$ K and $20$ K, for planets forming further away from the star.
-We note that this is not the only parameter degeneracy; most notably, the unavailability of nearby gas and dust can also cause a protoplanet to stop accreting, and the migration of a planet away from the star can likewise change its accretion rate.
-Therefore, while one can interpret the observation of a large gas giant as a constraint on DM, the non-observation cannot directly be interpreted as positive evidence.
+In Fig.
+1 we show the DM fraction needed to halt the formation process for Jupiter in the solar system, using $\rho_{\rm DM}=0.42\,\rm GeV\,cm^{-3}$ and $v_{\rm DM}=270\rm km\,s^{-1}$, we find that $f_{\rm cap}$ increases faster than the fraction necessary to satisfy (3) as a function of time.
+This does not go on indefinitely, because the rate of gas accumulation rises with increasing mass.
+Formation studies have shown that after the planet has grown to about $30M_{\oplus}$ gas capture rises exponentially, however as is seen for the benchmark in Fig.
+1 the annihilation of DM captured by the envelope heats the protoplanet enough to halt the formation at about $1.1\,\rm Myr$, when it is approximately $20\rm M_{\oplus}$.
+In Figs.2 and 3 we show constraints on the spin-dependent and spin-independent scattering cross sections respectively from the fact that the formation of Jupiter in our solar system has not stopped at a mass of $30\rm M_{\oplus}$ as purple contours.
 For comparison we show current direct detection constraints as gray contours.
 Here we show the maximal reach in the sensitivity to the direct nuclear recoil process, which are least model dependent.
-The reach is dominated by the CRESST results at lower masses for spin-independent Angloher et al.
-(2023), and spin-dependent Angloher et al.
-(2022) cross sections, while at higher masses the LUX Akerib et al.
-(2017) results dominate for spin-dependent interactions, while the XENON experiment dominates the spin-independent sensitivity Aprile et al.
-(2023).
-For large spin-independent cross sections, we also show the XQC constraint Erickcek et al.
-(2007); Mahdawi and Farrar (2018).
-Note that searches based on the Migdal effect Zhang et al.
-(2022) could potentially reach lower DM masses, however, the results have large theoretical uncertainties Cox et al.
-(2023), and are not displayed here.
+The reach is dominated by the CRESST results at lower masses for spin-independent CRESST:2022lqw, and spin-dependent CRESST:2022dtl cross sections, while at higher masses the LUX LUX:2017ree results dominate for spin-dependent interactions, while the XENON experiment dominates the spin-independent sensitivity XENONCollaboration:2023orw.
+For large spin-independent cross sections, we also show the XQC constraint Erickcek:2007jv; Mahdawi:2018euy.
+Note that searches based on the Migdal effect CDEX:2022kcd could potentially reach lower DM masses, however, the results have large theoretical uncertainties Cox:2022ekg, and are not displayed here.
 The sensitivity ceiling of direct detection experiments at large cross sections for light DM from Ref.
-Cappiello (2023) and for heavy DM from Refs.
-Kavanagh (2018); Cappiello et al.
-(2021) is shown.
-Since at the large cross sections considered the scaling with atomic mass number is uncertain Digman et al.
-(2019), we do not rescale the results for spin-dependent interactions and a more detailed analysis to obtain a precise ceiling values is needed.
-Note that in contrast the sensitivity of our constraints and suggested search is only affected at cross sections above $10^{-22}\text{cm}^{2}$, where large drift times would require a more detailed analysis of the DM distribution Acevedo et al.
-(2021); Leane and Smirnov (2022) and annihilation outside the core.
-We also show complementary cosmology-dependent constraints from the CMB Boddy and Gluscevic (2018).
-DM which scatters with a velocity-independent cross-section can also be constrained from the lack of MW satellites produced in such models Nadler et al.
-(2019).
-However, as the DM scattering rates could be different in the early universe Elor et al.
-(2023), those bounds are not as universal as the shown late-time sensitivity.
+Cappiello:2023hza and for heavy DM from Refs.
+Kavanagh:2017cru; Cappiello:2020lbk is shown.
+Since at the large cross sections considered the scaling with atomic mass number is uncertain Digman:2019wdm, we do not resale the results for spin-dependent interactions and a more detailed analysis to obtain a precise ceiling values is needed.
+Note that in contrast the sensitivity of our constraints and suggested search is only affected at cross sections above $10^{-22}\text{cm}^{2}$, where large drift times would require a more detailed analysis of the DM distribution Acevedo:2020gro; Leane:2022hkk and annihilation outside the core.
+We do not show complementary cosmology-dependent constraints such as those from the CMB and Lymann-$\alpha$.
+DM which scatters with a velocity-independent cross-section can also be constrained from the lack of MW satellites produced in such models Nadler:2019zrb.
+However, as the DM scattering rates could be different in the early universe Elor:2021swj, those bounds are not as universal as the shown late-time sensitivity.
+Finally, in the case of DM that comprises only a sub-fraction of the total DM abundance, the direct detection constraints are essentially insensitive below the GeV mass scale McKeen:2022poo (and the sensitivity of other constraints also shrinks significantly Li:2022idr), while our search covers significant parameter space even at a DM fraction as low as $f_{\rm DM}\sim 10^{-3}$, shown as solid chartreuse lines.
+In Figs.2 and 3, we show the sensitivity to DM interactions from observations of Jovians near the GC, for which we have assumed $\rho_{\rm DM}=100\,\rm GeV\,cm^{-3}$ and $v_{\rm DM}=10\rm\,km\,s^{-1}$, where at small cross sections we find that capture by the core is more efficient than capture by the envelope in both the spin-dependent and the spin-independent case.
+During the period of formation we are interested in, the core-grow is not appreciable, such that $f_{\rm cap}$ remains approximately constant.
+As the temperature increase needed to satisfy (2) does grow as a function of time, we expect that formation of Jovian planets in the GC halts at $10\rm M_{\oplus}$ in the parameter range indicated by the dashed contours, and proceeds as normal below these contours.
+Note that the premature ending is already in tension with inferences of Jovian planets in the GC from lensing events Sahu:2006ex; MOA:2009dkq; koshimoto2021no.
+Figure 4: Benchmark for spin-dependent dark matter: $m_{\chi}=5$ GeV, $\sigma_{p}=2\times 10^{-28}\,\rm cm^{2}$.
+The continuous lines show the capture rates at the time when the planet is $10\,\rm M_{\oplus}$ and the dashed lines show the rates at $30\,\rm M_{\oplus}$.
+In Fig.
+4 we show the capture efficiency as a function of distance from the GC assuming an NFW profile and a Maxwellian velocity dispersion.
+For small DM scattering cross sections, core capture dominates, preventing any gas accumulation for Jovian planets at some distance from the GC.
+For larger DM cross sections, envelope capture dominates, leading to the formation of Jovian planets with a decreasing mass the closer to the GC they are formed.
+This distinct phenomenology can be targeted by future observations of exoplanets.
 Lastly, as a result of DM heat injection, the proto-planet will respond by radiating away more energy.
@@
 Discussion
-In this paper, we have studied the inhibition of the formation of Jupiter in our solar system and Jovian planets in the Milky Way galaxy by the capture and annihilation of DM.
+In this letter, we have studied the inhibition of the formation of Jupiter in our solar system and Jovian planets in the Milky Way galaxy by the capture and annihilation of DM.
 We have shown that the heat generated by DM annihilation can increase the temperature of the protoplanet above the temperature at which the gas mixture evaporates, and determined the corresponding constraints on the parameter space of spin-(in)dependent DM.
 These constraints are competitive for relatively strongly interacting DM candidates, above the “ceiling” of direct detection constraints.
-We have shown that these constraints depend on the ambient temperature at formation, with stronger effects for planets forming further away from the star.
+We have shown that these constraints are not significantly weakened if only a subfraction of DM is responsible for the heating of protoplanets.
 We have also estimated the projected sensitivities from the existence of Jovian planets in the GC, where the DM density is higher and the typical velocity dispersion lower.
 Future expolanet observations by JWST and Roman may be used to test these scenarios.
-Moreover, as we explain in section Implications, for a particular DM (sub-component) mass and cross section the formation of Jovian planets is affected as a function of the distance from the GC, enabling a differential detection of the heating signal.
-In this work, we have estimated the sensitivity of protoplanets to DM injections using a semi-analytical calculation, with input from simulations in the absence of these injections.
-Future work should focus on the embedding of DM heat in planetary formation simulations, such that more realistic projections can be obtained.
+Moreover, as we explained in the previous section, for a particular DM (sub-component) mass and cross section the formation of Jovian planets is affected as a function of the distance from the GC, enabling a differential detection of the heating signal.
 Here we have focused on the core-accretion gas-capture theory of giant planet formation.
@@
 However, unlike hydrogen, as we demonstrate below the DM particles are concentrated in the core of the objects.
-Thus, in the short mean free path regime the lower number density of DM particles in the envelope implies a lower flux of escaping particles (Disrupting Jovian Formation).
+Thus, in the short mean free path regime the lower number density of DM particles in the envelope implies a lower flux of escaping particles (2).
 Here we estimate the evaporation rate of DM particles of a given mass, and compare it to the capture rate.
@@
 (11) form Ref.
-Leane and Smirnov (2022), under the assumption that the envelope has constant density $\rho_{E}$ (at $t=\rm Myr$, $\sim 3\times 10^{-8}\rm g\,cm^{-3}$) and the core has a constant density $\rho_{\rm core}$ (at $t=\rm Myr$, $\sim 2\rm g\,cm^{-3}$).
+Leane:2022hkk, under the assumption that the envelope has constant density $\rho_{E}$ (at $t=\rm Myr$, $\sim 3\times 10^{-8}\rm g\,cm^{-3}$) and the core has a constant density $\rho_{\rm core}$ (at $t=\rm Myr$, $\sim 2\rm g\,cm^{-3}$).
 We distinguish two regimes, the isothermal regime, where the DM mean free path $\lambda$ in the object is of the order of the object, and DM thermalises with the core at a constant temperature.
@@
 $g(r)=\theta\left(r-\frac{R_{\rm core}}{R_{E}}\right)\left(\frac{T_{E}}{T_{\rm core}}\right){}^{\frac{1}{2}\left(\frac{m_{\chi}}{m_{\text{p}}}+1\right)^{-3/2}-1}$
-$+\theta\left(\frac{R_{\rm core}}{R_{E}}-r\right)\,,$ (19)
-where $\theta$ indicates the Heaviside function.
-For the regime where the mean free path of DM becomes comparable to SM particles the above equation provides us with the DM density within the last scattering surface, to which we apply the Jeans’ escape criterion, demanding that the total surface evaporation rate is below the capture rate.
+$+\theta\left(\frac{R_{\rm core}}{R_{E}}-r\right)\,.$ (13)
+For the LTE regime the above equation provides us with the DM density within the last scattering surface, to which we apply the Jeans’ escape criterion, demanding that the total surface evaporation rate is below the capture rate.
 While for the isothermal, and transition regime we use the formalism of Ref.
-Gould (1987, 1990) to determine the evaporation mass comparing the evaporation rate and the time-scale needed to reach annihilation equilibrium assuming a core temperature of $3000$ K.
+1987ApJ...321..560G; 1990ApJ...356..302G, which has been applied to various celestail bodies in Ref.
+Garani:2021feo, to determine the evaporation mass assuming a core temperature of $3000$ K.
 This leads to the evaporation contour, shown by the dotted lines in Figs.
@@
 Note that this line is only marking to lowest mass of the DM retained in the object if DM interacts through contact interactions only.
-In the presence of attractive long range forces, the evaporation contours are significantly moved to lower DM masses, as discussed in Ref.
-Acevedo et al.
-(2023).
+In the presence of long range attractive forces, the evaporation contours are significantly moved to lower DM masses, as discussed in Ref.
+Acevedo:2023owd.